OpenStudy (anonymous):

4 subtracted from x is equal to x divided by 3. What is x?

OpenStudy (yttrium):

x-4 = x/3 Solve for x. :))

OpenStudy (ashy98):

x - 4 = x/3 multiply both sides by 3 3x - 12 = x add -x to both sides 2x - 12 = 0 add 12 to both sides 2x = 12 divide both sides by 2 and what do you get ???

OpenStudy (anonymous):

x=6 ?

OpenStudy (ashy98):

yes !(:

OpenStudy (anonymous):

thank you for helpin guys teehee!(:

OpenStudy (ashy98):

your welcome (: